# Accelerated creation of NOON states with ultracold atoms via counterdiabatic driving

### v3

@@ Abstract
A quantum control protocol is proposed for the creation of NOON states with $N$ ultracold bosonic atoms on two modes, corresponding to the coherent superposition $|N,0\rangle+|0,N\rangle$. This state can be prepared by using a third mode where all bosons are initially placed and which is symmetrically coupled to the two other modes. Tuning the energy of this third mode across the energy level of the other modes allows the adiabatic creation of the NOON state. While this process normally takes too much time to be of practical usefulness, due to the smallness of the involved spectral gap, it can be drastically boosted through counterdiabatic driving which allows for efficient gap engineering. We demonstrate that this process can be implemented in terms of static parameter adaptations that are experimentally feasible with ultracold quantum gases. Gain factors in the required protocol speed are obtained that increase exponentially with the number of involved atoms and thus counterbalance the exponentially slow collective tunneling process underlying this adiabatic transition.
@@
The concept of entanglement is of fundamental importance in quantum physics [1] and lies at the heart of various protocols in quantum information science [2]. A particularly intriguing category of entangled states are NOON states, given by a coherent superposition of the form $|N,0\rangle+e^{i\varphi}|0,N\rangle$ that involves $N$ bosons distributed over two modes. They can be seen as microscopic Schrödinger cat states and are of great interest in the context of quantum metrology [3, 4, 5, 6, 7] especially when being realized with massive bosons.
@@
This protocol is optimally executed in the self-trapping parameter regime [25] which is characterized by an onsite interaction strength $U$ that is large compared to the tunneling rate $J$ between the wells. In this regime, the many-body spectrum is energetically divided into multiple groups of levels. In particular, states where all particles are in the same well are all located at the top of the spectrum, isolated from the rest by a gap of size $\simeq U(N-1)/J$. The presence of interaction is thus of crucial importance for our method, as it induces a gap-protected reduced system that is useful for engineering the NOON state creation.
Figure 2: (a) Spectrum of the Hamiltonian as a function of the driving parameter $\varepsilon$, for $U=20J$ and $N=4$, focusing on the three relevant levels that are mainly composed of $|N,0,0\rangle,|0,N,0\rangle$ and $|0,0,N\rangle$. The red curves correspond to the original driven Hamiltonian (G method), while the blue curves represent the modified spectrum obtained through our GCD method for $T=10^{3}\hbar/J$, which combines an optimized driving protocol with the counterdiabatic Hamiltonian. Antisymmetric states are depicted with dashed lines. (b) Detection probabilities for a fixed total protocol time $T=3\times 10^{3}\hbar/J$. Thick curves indicates the probability of finding the system in the state $|\text{NOON}\rangle=(|N,0,0\rangle+|0,0,N\rangle)/\sqrt{2}$, while thin curves relates to the probability of measuring the state $|0,N,0\rangle$. One can observe a nearly perfect transition with the GCD method, in contrast to method G, which only attains about 10% of population inversion.
@@
$g^{(n)}=\sum_{m\neq n}\frac{\innerproduct{n|\partial_{\varepsilon}H}{m}\innerproduct{m|\partial_{\varepsilon}H}{n}}{(E_{n}-E_{m})^{2}}$ (6)
@@
A remaining challenge is that the counterdiabatic Hamiltonian (9) is defined within the reduced 3-level system spanned by the states $|N,0,0\rangle,|0,N,0\rangle$ and $|0,0,N\rangle$ (or, more precisely, by the perturbative modifications thereof, see [40]) but need, in practice, to be implemented in the framework of the Bose-Hubbard system under consideration. Similarly as in [58, 59, 60], we propose to emulate the requested effective Hamiltonian $H_{\text{red}}+H_{\text{CD}}$ via a suitable modification of the physical parameters $U,J$ and $\varepsilon$ of our system, thus incorporating the action of the counterdiabatic Hamiltonian into experimentally controllable parameters. We therefore define effective parameters $U_{\text{eff}},J_{\text{eff}}$ and $\varepsilon_{\text{eff}}(t)$ such that
@@
$\varepsilon_{\text{eff}}(t)=\varepsilon(t)+[\tilde{\mathcal{E}}(U,J)-\tilde{\mathcal{E}}(U_{\text{eff}},J_{\text{eff}})]/\mathcal{N},$ (14)
where
@@
As previously mentioned, the combination of counterdiabatic evolution with geodesic driving allows us to define $\Omega$ as time-independent. Consequently, the effective parameters $U_{\text{eff}}$ and $J_{\text{eff}}$ are also time-independent and can be properly tuned at the beginning of the protocol, and the only temporal dependence that needs to be managed is the driving of the energy of the central site. In optical lattices, this can be achieved by modulating the frequency difference between the lasers defining the lattice [61, 62, 63, 40].
@@
Results. Figure 3 displays the numerically computed infidelity as a function of the total protocol time $JT/\hbar$ for $N=4$ and different values of $U/J$. This infidelity is defined as $1-F$ with $F=|\langle\text{NOON}|\psi(T)\rangle|^{2}$ and $|\text{NOON}\rangle=(|N,0,0\rangle+|0,0,N\rangle)/\sqrt{2}$, where $|\psi(t)\rangle$ is the time-dependent state vector of the system starting at $|\psi(0)\rangle=|0,N,0\rangle$. To quantify the impact of counterdiabatic driving, we define by $T_{G}$ and $T_{GCD}$ the protocol time needed to reach the fidelity $F=2^{-1/2}=0.71$ with the geodesic protocol (G) and the combined geodesic and counterdiabatic driving (GCD), respectively. For $N=4$ and $U/J=20$, a protocol speed gain factor $g=T_{G}/T_{GCD}\simeq 9.5\times 10^{3}$ is calculated. Important gains in the protocol time are also obtained for high fidelities, such as $F=0.99$, as shown in Fig. 3.
@@
Figure 4: (a),(b) Gain ratio between the times required to achieve a fidelity of $F=0.71$ using the G and GCD methods for various interaction strengths, as a function of the number of particles. We find an exponential increase in efficiency $T_{G}\sim(J/U)^{N-1}$ when using the GCD method compared to the G method. (c),(d),(e) Protocol time $JT/\hbar$ as a function of particle number $N$ needed to yield the fidelity (a) $F=0.7$, (b) $F=0.9$, (c) $F=0.99$ at fixed $NU/J=60$. The exponential increase of $T$ with $N$, given by $T_{G}\sim(J/U)^{N-1}$ in absence of counterdiabatic driving, is substantially amended by the GCD method.
To better understand these results in terms of a specific physical context, let us consider a gas of 87Rb atoms, with $m=1.443\times 10^{-25}$ kg and the s-wave scattering length $a_{s}=5.313$ nm, which is confined in an optical lattice that is generated by lasers with the wavelength $\lambda=1064$ nm, whose lattice site energies are tuned to achieve the adiabatic transition (see [40] for a more detailed description of a possible experimental realization). According to [64, 65, 66, 26], a characteristic hopping time scale in case of the ratio $U/J=20$ is then given by $\hbar/J=4.4\times 10^{-3}$ s. Using this time scale for the example case depicted in Fig. 3, a fidelity of $F=0.98$ is obtained after a duration of $0.26$ s with the GCD method. In contrast, using only a geodesic driving will require at least $97$ s.
Conclusion. We demonstrated the feasibility of generating stable NOON states using advanced adiabatic techniques. Specifically, the creation of NOON states can be achieved through the use of the geodesics of the parameter manifold, which is to be complemented by the addition of a time-independent counterdiabatic Hamiltonian. This latter key ingredient allows for efficient gap engineering and gives rise to a drastic reduction of the creation time as compared to a purely adiabatic transition protocol [18], while maintaining its inherent robustness and a highly satisfactory purity of the NOON state.
@@ Acknowledgements.
for a complex hopping. Let us denote the interaction energy shifted by the driving term $\varepsilon$ as $E_{n_{1},n_{2},n_{3}}=n_{2}\varepsilon+\sum_{i=1}^{3}Un_{i}(n_{i}-1)/2$ and the hopping terms as $\mathcal{J}_{n_{i},n_{2}}=J\sqrt{n_{i}(n_{2}+1)}$ and $\mathcal{J}_{n_{2},n_{i}}^{*}=J^{*}\sqrt{n_{2}(n_{i}+1)}$, where $n_{i}$ are the number of particles in site $i$. Using the stationary Schrödinger equation, we end up with a system of coupled equations
[…]
@@
$\mathcal{E}(U,J)=\frac{\mathcal{J}_{N\to 0}^{2}}{E^{(2)}-\frac{\mathcal{J}_{N-1,1,0}^{2}}{E-E_{N-2,2,0}}-\frac{\mathcal{J}_{1\to 0}^{2}}{E-E_{N-1,0,1}}}$ (9)
where $E^{(2)}=E_{N,0,0}-E_{N-1,1,0}+\mathcal{J}_{N\to 0}^{2}/(E_{N,0,0}-E_{N-1,1,0})$. Expanding for $E\gg\mathcal{J}$, the corrective term is
$\mathcal{E}(U,J)=\frac{NJ^{2}}{(N-1)U}+\frac{NJ^{4}}{(N-1)^{2}(N-2)U^{3}}$ (10)
@@
$+\frac{N|J|^{2}}{(N-1)U}\varepsilon+\mathcal{N}\varepsilon$
@@
$+\frac{N|J|^{2}}{(N-1)U}\varepsilon$
We thus obtain the following expression for the shifted number of particles:
$\mathcal{N}=N+\delta N=N[1-3|J|^{2}/((N-1)U)^{2}].$ (14)
The effective coupling term between $|N,0,0\rangle$ and $|0,N,0\rangle$ is calculated from the trajectory depicted in bold in Fig. 1. By symmetry, the coupling must be the same between $|0,N,0\rangle$ and $|0,0,N\rangle$. We obtain
@@
At time $t=0$, a quench is to be implemented by which means the on-site energies of the populated lattice sites becomes suddenly increased, thus slightly exceeding the energies on the adjacent sites. As illustrated in Fig. 2(b), this can be achieved by suddenly increasing the amplitude $\eta$ from $-4$ to 00. The resulting (two-period) superlattice configuration can be considered to be the starting point for the NOON state creation process to be implemented. The amplitude $\eta$ has then to be slowly decreased, according to the protocol that we discuss in the main text of the paper. In addition, an artificial gauge field has to be induced along the lattice [1, 2] in order to give rise to the complex hopping matrix element (see Eq. (10) of the manuscript) that is required for the implementation of the counterdiabatic driving protocol. Additional (even time-dependent) adaptations of the Bose-Hubbard parameters can, if needed, be achieved by suitable variations of the global lattice amplitude $V_{0}$ as well as of the analogous amplitudes along the other two axes of the lattice.

### v4

@@ Abstract
A quantum control protocol is proposed for the creation of NOON states with $N$ ultracold bosonic atoms on two modes, corresponding to the coherent superposition $|N,0\rangle+|0,N\rangle$. This state can be prepared by using a third mode where all bosons are initially placed and which is symmetrically coupled to the two other modes. Tuning the energy of this third mode across the energy level of the other modes allows the adiabatic creation of the NOON state. While this process normally takes too much time to be of practical usefulness, due to the smallness of the involved spectral gap, it can be drastically boosted through counterdiabatic driving which allows for efficient gap engineering. We demonstrate that this process can be implemented in terms of static parameter adaptations that are experimentally feasible with ultracold quantum gases. Gain factors in the required protocol speed are obtained that increase exponentially with the number of involved atoms and thus counterbalance the exponentially slow collective tunneling process underlying this adiabatic transition. Besides optimizing the protocol speed, our NOON-state preparation scheme achieves excellent fidelities that are competitive for practical applications.
@@
The concept of entanglement is of fundamental importance in quantum physics [1] and lies at the heart of various protocols in quantum information science [2]. A particularly intriguing category of entangled states are NOON states, given by a coherent superposition of the form $|N,0\rangle+e^{i\varphi}|0,N\rangle$ that involves $N$ bosons distributed over two modes. They can be seen as microscopic Schrödinger cat states and are of great interest in the context of quantum metrology [3, 4, 5, 6, 7, 8] especially when being realized with massive bosons.
@@
This protocol is optimally executed in the self-trapping parameter regime [26] which is characterized by an onsite interaction strength $U$ that is large compared to the tunneling rate $J$ between the wells. In this regime, the many-body spectrum is energetically divided into multiple groups of levels. In particular, states where all particles are in the same well are all located at the top of the spectrum, isolated from the rest by a gap of size $\simeq U(N-1)/J$ (see Fig. 2 (a)). The presence of interaction is thus of crucial importance for our method, as it induces a gap-protected reduced system that is useful for engineering the NOON state creation.
Figure 2: (a) Spectrum of the Hamiltonian (1) as a function of the driving parameter $\varepsilon$, for $U=20J$ and $N=4$, with a geodesic driving $G$. The gap region is depicted by a black rectangle. (b) Focus on the three relevant levels of the spectrum (a) that are mainly composed of $|N,0,0\rangle,|0,N,0\rangle$ and $|0,0,N\rangle$. The red curves correspond to the original driven Hamiltonian (G method), while the blue curves represent the modified spectrum obtained through our GCD method for $T=3\times 10^{3}\hbar/J$, which combines an optimized driving protocol with the counterdiabatic Hamiltonian. Antisymmetric states are depicted with dashed lines. (c) Detection probabilities for a fixed total protocol time $T=3\times 10^{3}\hbar/J$. Thick curves indicates the probability of finding the system in the state $|\text{NOON}\rangle=(|N,0,0\rangle+|0,0,N\rangle)/\sqrt{2}$, while thin curves relates to the probability of measuring the state $|0,N,0\rangle$. One can observe a nearly perfect transition with the GCD method, in contrast to method G, which only attains about 10% of population inversion.
@@
$g^{(n)}=\text{Re}\sum_{m\neq n}\frac{\innerproduct{n|\partial_{\varepsilon}H}{m}\innerproduct{m|\partial_{\varepsilon}H}{n}}{(E_{n}-E_{m})^{2}}$ (6)
@@
A remaining challenge is that the counterdiabatic Hamiltonian (9) is defined within the reduced 3-level system spanned by the perturbative modifications of the states $|N,0,0\rangle,|0,N,0\rangle$ and $|0,0,N\rangle$ (see [41]). In practice, however, it needs to be implemented in the framework of the Bose-Hubbard system under consideration. Similarly as in [59, 60, 61], we propose to emulate the requested effective Hamiltonian $H_{\text{red}}+H_{\text{CD}}$ via a suitable modification of the physical parameters $U,J$ and $\varepsilon$ of our system, thus incorporating the action of the counterdiabatic Hamiltonian into experimentally controllable parameters. We therefore define effective parameters $U_{\text{eff}},J_{\text{eff}}$ and $\varepsilon_{\text{eff}}(t)$ such that
@@
$\varepsilon_{\text{eff}}(t)=\mathcal{N}\varepsilon(t)/\mathcal{N}_{\text{eff}}+[\tilde{\mathcal{E}}(U,J)-\tilde{\mathcal{E}}(U_{\text{eff}},J_{\text{eff}})]/\mathcal{N}_{\text{eff}},$ (14)
with $\mathcal{N}_{\text{eff}}=\mathcal{N}(U_{\text{eff}},J_{\text{eff}})$ and where
@@
As previously mentioned, the combination of counterdiabatic evolution with geodesic driving allows us to define $\Omega$ as time-independent. Consequently, the effective parameters $U_{\text{eff}}$ and $J_{\text{eff}}$ are also time-independent and can be properly tuned at the beginning of the protocol, and the only temporal dependence that needs to be managed is the driving of the energy of the central site. In optical lattices, this can be achieved by modulating the frequency difference between the lasers defining the lattice (see Supplemental Material [41] and references [62, 63, 64]).
@@
Results. Figure 3 displays the numerically computed infidelity as a function of the total protocol time $JT/\hbar$ for $N=4$ and different values of $U/J$. This infidelity is defined as $1-F$ with $F=|\langle\text{NOON}|\psi(T)\rangle|^{2}$ and $|\text{NOON}\rangle=(|N,0,0\rangle+|0,0,N\rangle)/\sqrt{2}$, where $|\psi(t)\rangle$ is the time-dependent state vector of the system starting at $|\psi(0)\rangle=|0,N,0\rangle$. To quantify the impact of counterdiabatic driving, we define by $T_{G}$ and $T_{GCD}$ the protocol time needed to reach the fidelity $F=0.99$ with the geodesic protocol (G) and the combined geodesic and counterdiabatic driving (GCD), respectively. For $N=4$ and $U/J=20$, a protocol speed gain factor $g=T_{G}/T_{GCD}\simeq 10^{4}$ is calculated. Important gains in the protocol time are also obtained for high fidelities, such as $F=0.999$, as shown in Fig. 3.
@@
Figure 4: (a),(b) Gain ratio between the times required to achieve a fidelity of $F=0.97$ using the G and GCD methods for various interaction strengths, as a function of the number of particles. We find an exponential increase in efficiency $T_{G}/T_{GCD}\sim(U/J)^{N-1}$ when using the GCD method compared to the G method. (c),(d),(e) Protocol time $JT/\hbar$ as a function of particle number $N$ needed to yield the fidelity (a) $F=0.97$, (b) $F=0.99$, (c) $F=0.999$ at fixed $NU/J=60$. The exponential increase of $T$ with $N$, given by $T_{G}/T_{GCD}\sim(U/J)^{N-1}$ in absence of counterdiabatic driving, is substantially amended by the GCD method.
To better understand these results in terms of a specific physical context, let us consider a gas of 87Rb atoms, with $m=1.443\times 10^{-25}$ kg and the s-wave scattering length $a_{s}=5.313$ nm, which is confined in an optical lattice that is generated by lasers with the wavelength $\lambda=1064$ nm, whose lattice site energies are tuned to achieve the adiabatic transition (see [41] for a more detailed description of a possible experimental realization). According to [65, 66, 67, 27], a characteristic hopping time scale in case of the ratio $U/J=20$ is then given by $\hbar/J=4.4\times 10^{-3}$ s. Using this time scale for the example case depicted in Fig. 3, a fidelity of $F=0.99$ is obtained after a duration of $30$ ms with the GCD method. In contrast, using only a geodesic driving will require at least $97$ s.
Conclusion. We demonstrated the feasibility of generating stable NOON states using advanced adiabatic techniques. Specifically, the creation of NOON states can be achieved through the use of the geodesics of the parameter manifold, which is to be complemented by the addition of a time-independent counterdiabatic Hamiltonian. This latter key ingredient allows for efficient gap engineering and gives rise to a drastic reduction of the creation time as compared to a purely adiabatic transition protocol [19], while maintaining its inherent robustness and a highly satisfactory purity of the NOON state. In particular, for $N=4$ particles, a 99% pure NOON state can be obtained after only a few tens of milliseconds.
@@ Acknowledgements.
for a complex hopping hopping parameter $J$. Let us denote the interaction energy shifted by the driving term $\varepsilon$ as $E_{n_{1},n_{2},n_{3}}=n_{2}\varepsilon+\sum_{i=1}^{3}Un_{i}(n_{i}-1)/2$ and the hopping terms as $\mathcal{J}_{n_{i},n_{2}}=J\sqrt{n_{i}(n_{2}+1)}$ and $\mathcal{J}_{n_{2},n_{i}}^{*}=J^{*}\sqrt{n_{2}(n_{i}+1)}$, where $n_{i}$ are the number of particles in site $i$. Using the stationary Schrödinger equation, we end up with a system of coupled equations
$E\Psi_{N,0,0}=E_{N,0,0}\Psi_{N,0,0}-\mathcal{J}_{N,0}\Psi_{N-1,1,0}$ (2)
@@
$\mathcal{E}(U,J)=\frac{|\mathcal{J}_{N,0}|^{2}}{E^{(2)}-\frac{|\mathcal{J}_{N-1,1}|^{2}}{E-E_{N-2,2,0}}-\frac{|\mathcal{J}_{1,0}|^{2}}{E-E_{N-1,0,1}}}$ (9)
where $E^{(2)}=E_{N,0,0}-E_{N-1,1,0}+|\mathcal{J}_{N,0}|^{2}/(E_{N,0,0}-E_{N-1,1,0})$. Expanding for $E\gg|\mathcal{J}|^{2}$, the corrective term is
$\mathcal{E}(U,J)=\frac{N|J|^{2}}{(N-1)U}+\frac{N|J|^{4}}{(N-1)^{2}(N-2)U^{3}}$ (10)
@@
$-\frac{2N|J|^{2}}{(N-1)^{2}U^{2}}\varepsilon+\mathcal{N}\varepsilon$
@@
$+\frac{N|J|^{2}}{(N-1)^{2}U^{2}}\varepsilon$
We thus obtain the expression :
$\mathcal{N}=N+\delta N=N[1-3|J|^{2}/((N-1)U)^{2}]\equiv\mathcal{N}(U,J).$ (14)
for the shifted number of particles. The effective coupling term between $|N,0,0\rangle$ and $|0,N,0\rangle$ is calculated from the trajectory depicted in bold in Fig. 1. By symmetry, the coupling must be the same between $|0,N,0\rangle$ and $|0,0,N\rangle$. We obtain
@@
We can verify the validity of the approximation by comparing the spectrum of $H_{\text{red}}$, given by
$E_{0}=\mathcal{E},$ (16)
$E_{\pm}=\left(\tilde{\mathcal{E}}+\mathcal{E}+\mathcal{N}\varepsilon\pm\sqrt{8\mathcal{J}^{2}+(\tilde{\mathcal{E}}-\mathcal{E}+\mathcal{N}\varepsilon)^{2}}\right)/2,$ (17)
with the three highest energy levels of the Hamiltonian (1). Figure 2 (a) shows the complete spectrum of the Bose-Hubbard Hamiltonian for a particle number $N=10$. A large gap of width $U(N-1)/J$ isolates the three highest levels from the rest of the spectrum. Figure 2(b) shows a zoom on the avoided crossing between the levels of interest. The quality of the approximation (15),(16) for the energy levels is also illustrated in Fig. 2(c) where these levels are compared to the upper part of the Bose-Hubbard spectrum for various particle numbers.
Figure 2: (a) Spectrum of the Bose-Hubbard Hamiltonian (black lines) as a function of the drive parameter $\varepsilon/J$, for $U=20J$ and $N=10$. The upper edge of the spectrum is isolated from the rest of the levels by a gap of width $U(N-1)/J$. The levels of the reduced Hamiltonian are shown with red dotted lines. (b) Zoom on the crossing of the three highest levels. (c) Value of the three highest energies shifted by the interaction energy with $U=20J$, as a function of the number of particles, for the undriven ($\varepsilon=0$) Bose-Hubbard Hamiltonian (BHM, in black) and for the reduced Hamiltonian (Hred, in red).
[…]
Fig. 2 shows the impact of the counterdiabatic term on the spectrum of the Bose-Hubbard Hamiltonian. In particular, the use of a GCD protocol indeed allows for widening the gap, proportional to $J(J/U)^{N-1}$, and thus reduces the time required for adiabatic dynamics. One can also observe that the counterdiabatic term shifts the energies, since the spectrum is modified by the addition of the term ensuring adiabaticity. This effect is easily verified by examining the difference in value between the antisymmetric level of the G protocol, equal to $\mathcal{E}(U,J)$, and that of the GCD protocol, equal to $\mathcal{E}(U_{\text{eff}},J_{\text{eff}})$. The right column shows that the fidelity obtained with the GCD protocol is excellent.
@@
At time $t=0$, a quench is to be implemented by which means the on-site energies of the populated lattice sites becomes suddenly increased, thus slightly exceeding the energies on the adjacent sites. As illustrated in Fig. 4(b), this can be achieved by suddenly increasing the amplitude $\eta$ from $-4$ to 00. The resulting (two-period) superlattice configuration can be considered to be the starting point for the NOON state creation process to be implemented. The amplitude $\eta$ has then to be slowly decreased, according to the protocol that we discuss in the main text of the paper. In addition, an artificial gauge field has to be induced along the lattice [1, 2, 3] in order to give rise to the complex hopping matrix element (see Eq. (10) of the manuscript) that is required for the implementation of the counterdiabatic driving protocol. Additional (even time-dependent) adaptations of the Bose-Hubbard parameters can, if needed, be achieved by suitable variations of the global lattice amplitude $V_{0}$ as well as of the analogous amplitudes along the other two axes of the lattice.